# A Framework for Holistic KLD-based Waveform Design for Multi-User-Multi-Target ISAC Systems

### v2

@@ Abstract
This paper introduces a novel framework that leverages the Kullback-Leibler Divergence (KLD) metric to analyse and optimise performance trade-offs in integrated sensing and communication (ISAC) systems. We consider a multiple-input multiple-output (MIMO) base station that simultaneously serves communication user equipments (UEs) and detects multiple targets using a shared antenna deployment. We apply this framework to two widely used communication beamforming techniques, maximum ratio transmission (MRT) and zero-forcing (ZF), to assess their impact on the radar subsystem’s performance. Additionally, two optimisation problems are formulated: the first optimises the radar subsystem’s KLD under communication constraints, and the second focuses on communication waveform KLD optimisation with constraints on the radar KLD. These problems are solved using a projected gradient method with adaptive penalties for the radar waveforms and a gradient-assisted interior point method (IPM) for the communication waveforms. Through theoretical derivations and extensive simulations, we demonstrate that our KLD approach effectively characterises and optimises the performance trade-offs between sensing and communication in ISAC systems. The results show significant improvements in both radar detection and communication performance when compared to traditional MRT and ZF beamforming, and the identity covariance design for radar subsystems. These findings promote a more holistic design and optimisation of ISAC for next-generation wireless networks and demonstrate the advantages of KLD-based optimisation in balancing the performance of both sensing and communication.
@@
Integrated sensing and communication, multiple-input-multiple-output (MIMO), radar, zero-forcing, maximum ratio transmission, beamforming, optimisation, Kullback–Leibler divergence (KLD).
@@ I Introduction
The evolution of ISAC has been accelerated by advances in multi-antenna technologies, particularly massive Multiple-Input Multiple-Output (MIMO) systems. These systems allow for the generation of highly directional beams and support simultaneous services to a large number of users, thereby significantly enhancing both communication capacity and sensing accuracy [16]. The synergy between ISAC and massive MIMO presents a promising avenue for meeting the stringent requirements of future wireless networks. However, evaluating the performance of ISAC systems poses unique challenges due to their dual functionality. Traditionally, communication and sensing subsystems have been assessed using distinct metrics, with communication performance measured by achievable rate, outage probability, and bit error rate (BER), while the sensing performance is typically evaluated using metrics like estimation rate, detection probability, false alarm probability, and mean square error (MSE) [17]. This disparity complicates the holistic assessment and optimisation of ISAC systems.
@@ I-A Literature review
To address these challenges mentioned above, there has been a growing interest in developing unified performance measures that can simultaneously capture both sensing and communication capabilities. One such measure that has gained traction is the Kullback-Leibler divergence (KLD), also known as the relative entropy. KLD has a well-established history in information theory and has been extensively used in the analysis and design of sensing systems, particularly in the context of MIMO radars [18, 19, 20]. It quantifies the dissimilarity between two probability distributions, serving as a measure of the information gain achieved by using one distribution over another. While the application of KLD in sensing systems is well-established, its potential in characterising communication system performance remains open. Recent studies have begun to bridge this gap, demonstrating the utility of KLD in assessing the performance trade-offs between sensing and communication functionalities in ISAC systems. For instance, in our previous works [21, 22], we employed KLD to analyse performance in systems with separated antenna deployment, while in [23] we proposed a low-complexity unified objective function based on KLD, aimed at optimising network resources for both separated and shared deployment scenarios. Moreover, further insights about the achievable KLD trade-off in ISAC systems are explored in [24], and more designs have been discovered. Nonetheless, the literature lacks a comprehensive investigation and beamforming design for a generalised multi-user-multi-target (MUMT) ISAC system using the promising KLD.
Further efforts have explored various optimisation approaches for ISAC systems, each addressing specific aspects of system performance. For example, joint radar-communication beamforming optimisation based on the Cramér-Rao Bound (CRB) has been proposed to improve both sensing and communication performance [25]. This approach minimises the CRB of angle estimation for radar while maximising the signal-to-interference-plus-noise ratio (SINR) for communication, thus enhancing the accuracy of sensing and the quality of communication simultaneously. However, these methods often focus on optimising individual components rather than adopting a unified framework that captures the interplay between different functionalities. Beamforming optimisation has been a particular focus, with studies investigating partially connected hybrid beamforming designs to address the hardware complexity of fully digital systems [26]. While these approaches offer practical advantages in terms of implementation, they do not fully capture the nuanced trade-offs between sensing and communication performance, especially in deeply integrated scenarios. Waveform optimisation has also been extensively studied, with recent work proposing a symbol-level precoding scheme using a faster-than-Nyquist approach to enhance the spectral efficiency in bandwidth-constrained environments [27]. This approach focuses on communication waveform design while considering the radar functionality as a constraint, ensuring that the designed waveforms maintain certain radar performance metrics. However, this work does not consider a unified optimisation framework that could simultaneously address the needs of both subsystems.
@@ I-B Motivation and contribution
Despite these advances, current approaches often fall short of fully exploiting the potential of integrated systems. Most notably, there is a need for more comprehensive models that can accurately reflect the real-world operation of ISAC systems, particularly in scenarios where communication signals can be utilised for radar target detection alongside dedicated radar signals. Exploiting the interference from communication signals as a source of information for radar operations enhances radar detection capabilities, improving the system’s overall resource efficiency [14]. This dual-use of communication signals represents a significant departure from traditional approaches, which typically treat interference as a detrimental factor to be mitigated rather than leveraged[28, 29].
Toward these objectives and motivated by the fact that KLD is able to provide a unified measure for ISAC system holistically, this paper proposes a novel framework based on the Kullback-Leibler divergence (KLD) to characterise and optimise the sensing-communication performance trade-offs in shared antenna deployment ISAC systems. This work addresses several gaps in the existing literature by introducing a unified KLD-based performance measure that captures the synergies and trade-offs between sensing and communication in ISAC systems. By considering the shared deployment scenario and proposing specific optimisation techniques for both subsystem waveforms, we provide a comprehensive approach for system optimisation that reflects the true potential of ISAC systems in future wireless networks. Additionally, our analysis of different beamforming techniques within this framework offers valuable guidance for designing ISAC systems to effectively balance the needs of both subsystem functionalities. Our system model consists of a multiple-input-multiple-output (MIMO) base station (BS) that simultaneously serves communication user equipment (UEs) and detects multiple targets using shared resources. The key contributions of this work are:
  • Unified Performance Measure: We derive the Kullback-Leibler Divergence (KLD) specifically for the integrated operation of communication and radar systems within ISAC. This performance measure unifies the communication error rate with radar target detection capabilities into a single objective function. By deriving this KLD-based metric for communication and radar components, we optimise ISAC system design holistically, addressing the trade-offs and enhancing overall system performance.
  • Exploitation of Communication Interference: Our framework accounts for the fact that communication signals can also be exploited for radar target detection alongside dedicated radar signals, providing a more comprehensive and useful model of ISAC system operation. This dual-use approach enhances the radar performance by utilising communication interference as an additional source of information for target detection.
@@
  • KLD-based Optimised ISAC: We propose two novel optimisation techniques: radar and communication waveform KLD-based optimisation. Both maximise their respective KLD measures subject to the minimum KLD of both subsystems and power constraints. The non-convex radar optimisation employs a projected gradient method with an adaptive penalty, while the communication optimisation uses a gradient-assisted interior point method. These techniques efficiently handle ISAC waveform design’s complex, multi-dimensional nature, enabling balanced performance optimisation across both subsystems within system constraints.
  • Benchmark Analysis: We investigate two benchmark beamforming techniques for the communication subsystem, namely maximum ratio transmission (MRT) and zero-forcing (ZF), and analyse their impact on the overall system performance within our KLD-based framework. This analysis provides practical insights for optimising ISAC performance in shared deployment scenarios, considering the complexity-performance trade-off.
@@
Our results demonstrate that the proposed KLD-based optimisation techniques significantly enhance ISAC system performance compared to the MRT and ZF scenarios for the communication subsystem and the radar subsystem’s conventional identity covariance design. The radar waveform optimisation shows substantial improvements in both target detection and communication performance, while the communication waveform optimisation primarily benefits the communication subsystem with modest radar performance gains. Both techniques exhibit robust performance across varying SNR levels, with the radar optimisation demonstrating particularly stable computational efficiency. Notably, incorporating communication signals in radar detection significantly enhances overall system performance, especially in scenarios with limited dedicated radar resources. These findings provide valuable insights for designing efficient shared deployment ISAC systems that balance sensing and communication requirements.
@@ I-C Paper organisation
The paper is structured as follows. Section II presents the system model. Sections III and IV analyse the communication and radar systems, respectively, including the derivations of KLD for each system. In Section V, the first KLD-based optimisation technique considers the radar KLD sum under communication KLD constraint is investigated. Section VI introduces the second KLD-based optimisation technique based on communication KLD sum with radar KLD constraint. Section VII displays the numerical results, and finally, Section VIII concludes the work.
@@ II System Model
We consider an $N$ antenna MIMO-BS. The antennas are utilised for detecting a maximum number of ${T}$ targets and serving $K$ number of single-antenna communication UEs in the downlink direction in shared deployment manner. The total transmitted power available at BS is $P_{\mathrm{T}}$ which is utilised for both sensing and data communication duties. The power $P_{\mathrm{T}}$ allocated to the radar and communication subsystems are denoted as $P_{\mathrm{r}}$ and $P_{\mathrm{c}}$, where $P_{\mathrm{T}}=P_{\mathrm{c}}+P_{\mathrm{r}}$. As a starting point, non-optimised beamforming techniques such as MRT and ZF are employed at BS to precode the information of communication UEs [16]. These techniques serve as benchmarks for our subsequent optimised systems, which will be introduced and analysed later in Sections V and VI. The combined transmitted ISAC signal $\mathbf{x}_{l}\in\mathbb{C}^{N\times 1}$ will be as follows,
@@
where $\mathbf{s}_{\mathrm{c},l}\in\mathbb{C}^{K\times 1}$ is a vector of communication UEs symbols, $\mathbf{s}_{\mathrm{r}}\in\mathbb{C}^{T\times 1}$ is a vector of the baseband radar waveforms for the potential targets, $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$ represents the precoding matrix for the communication subsystem, and $\mathbf{W}_{\mathrm{r},l}\in\mathbb{C}^{N\times T}$ represents the precoding matrix for the radar subsystem. The transmitted signals vector $\mathbf{x}_{l}$ is transmitted from the BS and received at the UEs as well as the targets which reflect the received $\mathbf{x}_{l}$ back to the BS. Although the radar signal component received at the UEs might cause interference that cannot be eliminated by the users, the communication signal component received at the targets and reflected back to the BS carries information about the target. Therefore, by carefully designing the precoding matrices at the BS, the sensing information can be enhanced by exploiting the communication signal. This dual-use approach enhances the radar performance by utilising the communication signal as an additional source of information. The radar waveform design is based on the combined transmit signal covariance matrix for the $t$-th beam where the covariance matrix of the $t$-th transmit beam is $\mathbf{R}_{t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}$, $\mathbf{x}_{t,l}=\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}+\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}$, and $L$ is the total number of snapshots. This formulation allows us to incorporate the communication signals into the radar detection process. It should be noticed that the power for the radar is integrated into the precoding vector $\mathbf{w}_{\mathrm{r},t,l}$, where it is designed to have $\|\mathbf{w}_{\mathrm{r},t,l}\|_{F}^{2}=\frac{P_{\mathrm{r}}}{T}$, and $\mathbf{w}_{\mathrm{r},t,l}=\frac{P_{\mathrm{r}}}{NT}\>\mathbf{\tilde{w}}_{\mathrm{r},t,l}$, where $\mathbf{\tilde{w}}_{\mathrm{r},t,l}$ is the $t$th beam precoding vector, and .$\|.\|_{F}$ is the Frobenius norm.
@@ II-A Communication System
where $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{h}^{2}\right)$ represents the channel from the BS to the $k$-th UE, $\omega_{\mathrm{MN},k}=\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}$ represents the IUI on the $k$-th UE from the other UEs, $d_{\mathrm{c},k}^{-\zeta/2}$ is the channel pathloss from BS to the $k$-th UE with $d_{\mathrm{c},k}$ representing the distance from BS to the $k$-th UE, and $\zeta$ is the pathloss exponent $\eta_{k}=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}}+n_{k,l}$ is the interference-plus-noise term with ${n}_{k,l}\sim\mathcal{C}\mathcal{N}\left(0,\>\sigma_{n}^{2}\right)$ representing the additive white Gaussian noise (AWGN). The precoding vector for the communication system is $\mathbf{w}_{\mathrm{c},k}\in\mathbb{C}^{N\times 1}$, which is typically designed based on the given channel matrix from MIMO-BS to UEs $\mathbf{H}\in\mathbb{C}^{N\times K}=\left[\mathbf{h}_{1},..,\mathbf{h}_{k},..,\mathbf{h}_{K}\right]$, where the elements of $\mathbf{H}$ are independent and identically distributed (i.i.d) complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$. In the next two subsections, we revisit the conventional MRT and ZF designs which are typically used to precode the communication data symbols at the BS. These precoding schemes will serve as benchmarks for our designs.
II-A1 MRT beamforming
In MRT precoding, the amount of received power at the UEs is maximised without taking into account the imposed interference. This approach usually leads to acceptable performance at low SNR values but its performance deteriorates at large SNRs. The normalised MRT beamforming matrix is designed for all UEs based on the channel matrix as follows,
$\mathbf{W}_{\mathrm{c}}=\mathbf{P}\>\mathbf{H}^{*}/\|\mathbf{H}\|_{F}=\mathbf{P}\>[\mathbf{h}_{1}^{*},\ldots,\mathbf{h}_{K}^{*}]/\|\mathbf{H}\|_{F},$ (3)
where $\mathbf{P}\in\mathbb{C}^{K\times K}=\mathrm{diag}\left(\sqrt{p_{1}},\cdots,\sqrt{p_{K}}\right)$ is the power allocation matrix for the communication users, which is a diagonal matrix that controls the power allocated for each UE.
II-A2 ZF beamforming
On the contrary, ZF precoding eliminates inter-user interference (IUI) at the UEs regardless of the amount of the received power. This approach usually results in excellent performance at high SNR values. The normalised ZF beamforming matrix is designed also based on the channel matrix, and can be represented as follows,
@@ II-A2 ZF beamforming
where $\mathbf{\tilde{W}}_{\mathrm{c}}=\mathbf{H}^{H}\>\left(\mathbf{H}\>\mathbf{H}^{H}\right)^{-1}$ is the ZF beamformer before normalisation.

### v4

@@ Abstract
This paper introduces a novel framework aimed at designing integrated waveforms for robust integrated sensing and communication (ISAC) systems. The system model consists of a multiple-input multiple-output (MIMO) base station that simultaneously serves communication user equipments (UEs) and detects multiple targets using a shared-antenna deployment scenario. By leveraging Kullback-Leibler divergence (KLD) to holistically characterise both communication and sensing subsystems, three optimisation problems are formulated: (i) radar waveform KLD maximisation under communication constraints, (ii) communication waveform KLD maximisation subject to radar KLD requirements, and (iii) an integrated waveform KLD-based optimisation for ISAC that jointly balances both subsystems. The first two problems are solved using a projected gradient method with adaptive penalties for the radar waveforms and a gradient-assisted interior point method (IPM) for the communication waveforms. The third, integrated waveform optimisation approach adopts an alternating direction method of multipliers (ADMM) framework to unify radar and communication waveform designs into a single integrated optimisation, thereby strengthening the trade-off between sensing and communication objectives and achieving higher overall performance than the individual radar- or communication-only techniques. Unlike most existing ISAC waveform designs that regard communication signals solely as interference for sensing, the proposed framework utilises the holistic ISAC waveform—that is, the superimposed communication and sensing signals—to boost detection performance in the radar subsystem. Simulation results show significant improvements in both radar detection and communication reliability compared with conventional zero-forcing beamforming and identity-covariance radar baselines, demonstrating the promise of KLD-based waveform designs for next-generation ISAC networks.
@@
Integrated sensing and communication, multiple-input-multiple-output (MIMO), radar, zero-forcing, beamforming, optimisation, Kullback–Leibler divergence (KLD).
@@ I Introduction
The evolution of ISAC has been accelerated by advances in multi-antenna technologies, particularly massive Multiple-Input Multiple-Output (MIMO) systems, which significantly enhance both communication capacity and sensing accuracy [16]. However, evaluating the performance of ISAC systems poses unique challenges due to their dual functionality. Traditionally, communication and sensing subsystems have been assessed using distinct metrics—communication performance through achievable rate, outage probability, and bit error rate (BER), and sensing performance through estimation rate, detection probability, false alarm probability, and mean square error (MSE) [17]. This disparity complicates the holistic assessment and optimisation of ISAC systems. To this end, some unified performance measures have been introduced in the literature, such as the mutual information [18, 19, 20], and Kullback-Leibler divergence, also known as the relative entropy, which we recently introduced [21, 22, 23, 24].
@@ I-A Literature review
Further efforts have explored various optimisation approaches for ISAC systems, each addressing specific aspects of system performance. For example, joint radar-communication beamforming optimisation based on the Cramér-Rao Bound (CRB) has been proposed to improve both sensing and communication performance [25], where this approach minimises the CRB of angle estimation while maximising the signal-to-interference-plus-noise ratio (SINR) for communication. However, these methods often focus on optimising individual components rather than adopting a unified framework that captures the interplay between different functionalities. Beamforming optimisation has been a particular focus, with studies investigating partially connected hybrid designs to address hardware complexity issues [26]. While these approaches offer some practical advantages in terms of implementation, they fail to fully capture the nuanced trade-offs between sensing and communication performance in deeply integrated scenarios. Similarly, waveform optimisation has also been extensively studied, with recent work using faster-than-Nyquist approaches to enhance spectral efficiency while maintaining radar functionality as a constraint [27]. Recent studies have also investigated joint transmit and receive beamforming in full-duplex ISAC systems to manage interference and enhance overall performance; however, these approaches primarily emphasise interference suppression rather than exploitation, thereby neglecting potential performance gains achievable through direct integration of communication signals into radar sensing processes [28]. Despite these advances, existing work has not adequately addressed the need for a unified optimisation framework that simultaneously serves both subsystems’ requirements and balances their performance trade-offs effectively [29]. Furthermore, a recent comprehensive survey on interference management in ISAC systems highlights that most current strategies remain oriented toward interference mitigation or cancellation, reinforcing the perspective of communication signals as detrimental rather than beneficial [30]. Consequently, such designs not only overlook opportunities to leverage inherent signal interactions but also risk introducing excessive complexity, which is especially problematic when real-time processing is required in dynamic environments.
@@ I-B Motivation and contribution
Despite these advances, current approaches often fall short of exploiting the full potential of the integrated system by assuming sensing and communication are competing subsystems rather than considering a synergic design. Most notably, there is a need for more comprehensive models that accurately reflect the real-world operation of ISAC systems, particularly in scenarios where communication signals can be utilised for radar target detection alongside dedicated radar signals. Exploiting the interference from communication signals as a source of information for radar operations, as shown in our paper, enhances radar detection capabilities, improving the system’s overall resource efficiency. This dual use of communication signals represents a significant departure from traditional approaches, which typically treat interference as a detrimental factor to be mitigated rather than leveraged[31, 32].
Toward these objectives and motivated by the fact that KLD is able to provide a unified measure for ISAC systems holistically, this paper proposes a novel ISAC waveform design with the objective of maximising KLD for the whole system as a design criterion in a shared antenna deployment scenario. Additionally, this work investigates the trade-off between trade-offs between sensing and communication in ISAC systems utilising KLD through well-established relations between KLD on one side and the detection probability in radar via Stein’s Lemma and BER in communication systems from the other side [22]. Our system model consists of a MIMO BS that simultaneously serves communication user equipment (UEs) and detects multiple targets using shared resources, using a statistical radar model. To this end, the key contributions of this work are:
@@
  • Two Novel Separate Waveform Optimisation Designs: we present radar and communication waveform optimisation using the unified KLD measure to represent the achievable performance for both subsystems on one scale. Both formulated problems aim at maximising the KLD measure of a certain subsystem subject to the minimum KLD of the other subsystem and power constraints. The non-convex radar optimisation employs a projected gradient method with an adaptive penalty, while the communication optimisation uses a gradient-assisted interior point method. These techniques efficiently handle ISAC waveform design’s complex nature, enabling balanced performance optimisation across both subsystems within system constraints.
  • Integrated Waveform Optimisation for ISAC: Beyond the individual radar and communication waveform-oriented designs, an integrated waveform optimisation framework for ISAC is introduced to optimise both waveforms together. This integrated waveform approach leverages an alternating direction method of multipliers (ADMM) algorithm, leading to enhanced performance in terms of radar detection accuracy and communication reliability compared with the separate waveform methods. By capturing and exploiting the synergy between sensing and data transmission, the unified technique achieves higher overall performance in shared-antenna ISAC deployments.
  • Exploitation of Communication Interference: Our framework accounts for the fact that communication signals can also be exploited for radar target detection alongside dedicated radar signals, providing a more comprehensive and useful model of ISAC system operation. This dual-use approach enhances the radar performance by utilising communication interference as an additional source of information for target detection, which offers new insights into the underlying synergy between radar and communication waveforms and quantifies how their joint optimisation can enhance overall ISAC capabilities.
  • Comprehensive Complexity Assessment: We provide a thorough computational complexity analysis of our proposed ISAC algorithms, evaluating their scalability and real-time feasibility. This analysis is critical for determining practical implementation strategies in dynamic network environments where swift adaptation and efficient resource utilisation are essential.
@@
Our results demonstrate that the proposed KLD-based optimisation techniques significantly outperform ZF beamforming and the radar subsystem’s identity covariance baseline. The radar waveform KLD-based optimisation achieves substantial gains in both target detection and communication quality, whereas the communication waveform KLD-based optimisation primarily benefits the communication subsystem, with modest radar performance gains. Both separate waveform techniques exhibit robust performance across varying SNR levels, with the radar waveform design demonstrating notably stable computational efficiency. The integrated waveform optimisation KLD-based for ISAC approach unifies both radar and communication waveforms into a single integrated design, delivering even higher overall performance in terms of sensing and communication signals. The numerical results highlight how leveraging communication signals for radar detection significantly enhances overall system performance, especially in scenarios with constrained dedicated radar resources. These findings emphasise the practical significance of KLD-based designs for shared-antenna ISAC systems, provide important adaptability and insights for balancing sensing and communication requirements in future wireless networks.
@@ I-C Paper organisation
The paper is structured as follows. Section II presents the system model. Sections III and IV analyse the communication and radar systems, respectively, including the derivations of KLD for each system. In Section V, the radar waveform optimisation is investigated using the KLD. Section VI, the communication waveform optimisation is investigated using the KLD. Section VII shows the complexity analysis for both waveform optimisation techniques, Section VIII displays the numerical results, and finally, Section IX concludes the work.
@@ II System Model
We consider an $N$ antenna MIMO-BS. The antennas are utilised for detecting a maximum number of ${T}$ targets and serving $K$ number of single-antenna communication UEs in the downlink direction in shared deployment manner. The total transmitted power available at BS is $P_{\mathrm{T}}$ which is utilised for both sensing and data communication duties. The power $P_{\mathrm{T}}$ allocated to the radar and communication subsystems are denoted as $P_{\mathrm{r}}$ and $P_{\mathrm{c}}$, where $P_{\mathrm{T}}=P_{\mathrm{c}}+P_{\mathrm{r}}$. As a starting point, non-optimised ZF beamforming technique is employed at BS to precode the information of communication UEs [16]. These techniques serve as benchmarks for our subsequent optimised systems, which will be introduced and analysed later in Sections V and VI. The combined transmitted ISAC signal $\mathbf{x}_{l}\in\mathbb{C}^{N\times 1}$ at the $l$-th snapshot can be formulated as follows,
@@
where $l\in\{1,2,\dots,L\}$ represents the discrete time index, and $L$ is the total number of snapshots considered for sensing outcome, $\mathbf{s}_{\mathrm{c},l}\in\mathbb{C}^{K\times 1}$ is a vector of communication UEs symbols, $\mathbf{s}_{\mathrm{r}}\in\mathbb{C}^{T\times 1}$ is a vector of the baseband radar waveforms for the potential targets, $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$ represents the precoding matrix for the communication subsystem, and $\mathbf{W}_{\mathrm{r},l}\in\mathbb{C}^{N\times T}$ represents the precoding matrix for the radar subsystem. The transmitted signals vector $\mathbf{x}_{l}$ is transmitted from the BS and received at the UEs as well as the targets which reflect the received $\mathbf{x}_{l}$ back to the BS. Although the radar signal component received at the UEs might cause interference that cannot be eliminated by the users, the communication signal component received at the targets and reflected back to the BS carries information about the target. Therefore, by carefully designing the precoding matrices at the BS, the sensing information can be enhanced by exploiting the communication signal. This dual-use approach enhances the radar performance by utilising the communication signal as an additional source of information. The radar waveform design is based on the combined transmit signal covariance matrix for the $t$-th beam where the covariance matrix of the $t$-th transmit beam is $\mathbf{R}_{t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}$, $\mathbf{x}_{t,l}=\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}+\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}$. This formulation allows us to incorporate the communication signals into the radar detection process. It should be noticed that the power for the radar is integrated into the precoding vector $\mathbf{w}_{\mathrm{r},t,l}$, where it is designed to have $\|\mathbf{w}_{\mathrm{r},t,l}\|_{F}^{2}=\frac{P_{\mathrm{r}}}{T}$, and $\mathbf{w}_{\mathrm{r},t,l}=\frac{P_{\mathrm{r}}}{NT}\>\mathbf{\tilde{w}}_{\mathrm{r},t,l}$, where $\mathbf{\tilde{w}}_{\mathrm{r},t,l}$ is the $t$th beam precoding vector, and .$\|.\|_{F}$ is the Frobenius norm.
@@ II-A Communication System
where $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{h}^{2}\right)$ represents the channel from the BS to the $k$-th UE, $\omega_{\mathrm{MN},k}=\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}$ represents the IUI on the $k$-th UE from the other UEs, $d_{\mathrm{c},k}^{-\zeta/2}$ is the channel pathloss from BS to the $k$-th UE with $d_{\mathrm{c},k}$ representing the distance from BS to the $k$-th UE, and $\zeta$ is the pathloss exponent $\eta_{k}=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r},l}+n_{k,l}$ is the interference-plus-noise term with ${n}_{k,l}\sim\mathcal{C}\mathcal{N}\left(0,\>\sigma_{n}^{2}\right)$ representing the additive white Gaussian noise (AWGN). The precoding vector for the communication system is $\mathbf{w}_{\mathrm{c},k}\in\mathbb{C}^{N\times 1}$, which is typically designed based on the given channel matrix from MIMO-BS to UEs $\mathbf{H}\in\mathbb{C}^{N\times K}=\left[\mathbf{h}_{1},..,\mathbf{h}_{k},..,\mathbf{h}_{K}\right]$, where the elements of $\mathbf{H}$ are independent and identically distributed (i.i.d) complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$. ZF designs which is typically used to precode the communication data symbols at the BS. This precoding scheme will serve as a benchmark for our designs. The ZF precoding eliminates inter-user interference (IUI) at the UEs regardless of the amount of the received power. This approach usually results in excellent performance at high SNR values. The normalised ZF beamforming matrix is designed also based on the channel matrix, and can be represented as follows,
@@
where $\mathbf{\tilde{W}}_{\mathrm{c}}=\mathbf{H}^{H}\>\left(\mathbf{H}\>\mathbf{H}^{H}\right)^{-1}$is the ZF beamformer before normalisation, and $\mathbf{P}\in\mathbb{C}^{K\times K}=\mathrm{diag}\left(\sqrt{p_{1}},\cdots,\sqrt{p_{K}}\right)$ is the power allocation matrix for the communication users, which is a diagonal matrix that controls the power allocated for each UE.